OpenStudy (anonymous):

Give the 10th term in the geometric sequence. 2, 6, 18, 54, ... Answer 39,366 13,122 4,374 5,400

OpenStudy (anonymous):

@Shadowys C or D?

OpenStudy (anonymous):

gimme a sec

OpenStudy (anonymous):

The answer is A

OpenStudy (anonymous):

a(n) = a*r^(n-1) a(10) = 2*3^(9) ------- a(10) = 39,366

OpenStudy (anonymous):

@Brittni0605 it A

OpenStudy (anonymous):

Natalie's right.

OpenStudy (anonymous):

Yes, @natalie135 is correct